# A Scalable Game-theoretic Approach to Urban Evacuation Routing and Scheduling

### v2

@@ Abstract
Evacuation planning is an essential part of disaster management where the goal is to relocate people under imminent danger to safety. However, finding jointly optimal evacuation routes and schedule that minimizes the average evacuation time or evacuation completion time, is a computationally hard problem. As a result, large-scale evacuation routing and scheduling continues to be a challenge. In this paper, we present a game-theoretic approach to tackle this problem. We start by formulating a strategic routing and scheduling game, named the Evacuation Game: Routing and Scheduling (egres), where players choose their route and time of departure. We show that: (i) every instance of egres has at least one pure strategy Nash equilibrium, and (ii) an optimal outcome in an instance will always be an equilibrium in that instance. We then provide bounds on how bad an equilibrium can be compared to an optimal outcome. Additionally, we present a polynomial-time algorithm, the Sequential Action Algorithm (saa), for finding equilibria in a given instance under a special condition. We use Virginia Beach City in Virginia, and Harris County in Houston, Texas as study areas and construct two egres instances. Our results show that, by utilizing saa, we can efficiently find equilibria in these instances that have social objective close to the optimal value.
@@ I Introduction
Evacuation plans are designed to ensure the safety of people living in areas that are prone to natural or man-made disasters such as hurricanes, tsunamis, wildfires, and toxic chemical spills. Large-scale evacuations have been carried out during the past hurricane seasons in Florida, Texas, Louisiana, and Mississippi. For instance, about $2.5$ million people were ordered to evacuate from Florida before the landfall of Hurricane Ian (2022) [1, 2]. Other examples of hurricanes when such large scale evacuations were carried out include Ida (2021), Laura (2020), Irma & Harvey (2017), Ike & Gustav (2008), Katrina & Rita (2005), Ivan (2004). It is, therefore, necessary to have effective and efficient evacuation plans in place to ensure the sustainability of cities and communities. Such a plan needs to determine two major components: $(i)$ Evacuation Routes, recommending the evacuees which route to take, and $(ii)$ Evacuation Schedule, recommending the evacuees when to leave.
@@ II Related Works
The evacuation routing and/or scheduling problem have been approached in several ways in the literature. A review of existing works can be found in the survey paper [5]. Hamacher and Tjandra [6] formulated the problem as a dynamic network flow optimization problem. They introduced the idea of time-expanded graphs that contain a copy of each node and each edge in the road network at each timestep within a time horizon. The authors then solved the problem using mathematical optimization methods. However, their proposed method had very high computational cost. This led the researchers to design and propose several heuristic methods [7, 8, 9]. The drawback of these methods is that they only consider the routing problem, i.e., they either do not consider the scheduling problem, or make simplifying assumptions such as evacuees leave at a constant rate.
@@ III Problem Formulation
    • $R_{i}$ is the set of all possible simple paths from source $src_{i}$ to any safe node in ${\mathcal{S}}$.
@@
Proof of Lemma 1 is provided in Appendix A. Let $a_{-i}$ denote actions of all players excluding player $i$ in outcome $a$, i.e. $a_{-i}=(a_{1},...,a_{i-1},a_{i+1},...,a_{n})$. Then:
@@
To quantify the quality of equilibria, we define Price of Anarchy (PoA) and Price of Stability (PoS) as follows:
@@
Figure 1(a) (evacuation network) and 1(b) (utility table) shows an example instance of egres. Here, we have two players: $P_{0}$ and $P_{1}$, corresponding to the two source nodes 00 and $1$ in Figure 1(a). There is only one safe node ($A$) in this instance. Each player has $1$ evacuee under her. Figure 1(b) shows the possible actions of the two players. Actions of $P_{0}$ are shown in rows, and actions of $P_{1}$ in columns. For instance, the action $[0,2,A],[(0,1)]$ of $P_{0}$ means: $P_{0}$ has chosen the route $[0,2,A]$, and one evacuee leaves from node 00 at timestep 00. Each cell of the utility table in Figure 1(b) corresponds to an outcome. The first and second value in each cell are the utility values of player $P_{0},P_{1}$ respectively. These values are calculated according to Equation 1.
@@ IV-A Existence of Pure Strategy Nash Equilibrium
Corollary 1 tells us that an equilibrium can be as good as an optimal outcome. Corollary 2, on the other hand, has important implications about the computation of equilibria. We discuss this further in Section IV-C.
@@ V-A Quality of the Equilibria
Figure 7: Box-plot showing the ratio of social objective at equilibrium to the lower bound of optimal social objective, of the $30$ equilibrium outcomes calculated by saa, for our two problem instances. The closer the ratio value to $1$, the better.
@@
Since, saa provides us equilibria that have confluent routes, we use the optimal confluent routes and schedule as our baseline. We use the Mixed Integer Program (MIP) formulation provided by Islam and Chen et al. [4] to find the optimal confluent routes and schedule for our two problem instances. We solved the MIPs with the MIP solver Gurobi [22]. However, since finding an optimal outcome is NP-hard and also hard to approximate [4], the solver was not able to provide us with exact optimal solutions. The solver did provide us with lower bounds for the optimal objective. We denote this lower bound by $opt_{lb}$. Let the social objective (i.e. sum of the utility of all players) in an equilibrium outcome be denoted by $opt_{eq}$. We then look at the ratio $opt_{eq}/opt_{lb}$ to quantify the quality of the equilibria found by saa. For example, if the ratio is $2$ for an equilibrium outcome, then it means the ratio of the social objective at that equilibrium to the optimal social objective is at-most $2$. The closer the ratio value to $1$, the better.
@@ V-B Scalability of saa
MIP solve time (minutes) $28$ $>120$ $>120$
SAA avg. run time (minutes) $14$ $20$ $27$
@@
MIP solve time (minutes) $56$ $>120$ $>120$
SAA avg. run time (minutes) $34$ $36$ $37$

### v3

@@ Abstract
Evacuation planning is an essential part of disaster management where the goal is to relocate people under imminent danger to safety. However, finding jointly optimal evacuation routes and a schedule that minimizes the average evacuation time or evacuation completion time, is a computationally hard problem. As a result, large-scale evacuation routing and scheduling continues to be a challenge. In this paper, we present a game-theoretic approach to tackle this problem. We start by formulating a strategic routing and scheduling game, named the Evacuation Game: Routing and Scheduling (egres), where players choose their route and time of departure. We show that: (i) every instance of egres has at least one pure strategy Nash equilibrium, and (ii) an optimal outcome in an instance will always be an equilibrium in that instance. We then provide bounds on how bad an equilibrium can be compared to an optimal outcome. Additionally, we present a polynomial-time algorithm, the Sequential Action Algorithm (saa), for finding equilibria in a given instance under a special condition. We use Virginia Beach City in Virginia, and Harris County in Houston, Texas as study areas and construct two egres instances. Our results show that, by utilizing saa, we can efficiently find equilibria in these instances that have social objective close to the optimal value.
@@ I Introduction
Evacuation plans are designed to ensure the safety of people living in areas that are prone to natural or man-made disasters such as hurricanes, tsunamis, wildfires, and toxic chemical spills. Large-scale evacuations have been carried out during the past hurricane seasons in Florida, Texas, Louisiana, and Mississippi. For instance, about $2.5$ million people were ordered to evacuate from Florida before the landfall of Hurricane Ian (2022) [1, 2]. Other examples of hurricanes when such large scale evacuations were carried out include Helene & Milton (2024), Idalia (2023), Ida (2021), Laura (2020), Irma & Harvey (2017), Ike & Gustav (2008), Katrina & Rita (2005). It is, therefore, necessary to have effective and efficient evacuation plans in place to ensure the sustainability of cities and communities. Such a plan needs to determine two major components: $(i)$ Evacuation Routes, recommending the evacuees which route to take, and $(ii)$ Evacuation Schedule, recommending the evacuees when to leave.
@@ II Related Works
The evacuation routing and/or scheduling problem have been approached in several ways in the literature. A review of existing works can be found in the survey paper [5]. Hamacher and Tjandra [6] formulated the problem as a dynamic network flow optimization problem. They introduced the idea of time-expanded graphs that contain a copy of each node and each edge in the road network at each timestep within a time horizon. The authors then solved the problem using mathematical optimization methods. However, their proposed method had very high computational cost. This led the researchers to design and propose several heuristic methods [7, 8, 9, 10]. The drawback of these methods is that they only consider the routing problem, i.e., they either do not consider the scheduling problem, or make simplifying assumptions such as evacuees leave at a constant rate.
@@ III Problem Formulation
    • $R_{i}$ is the set of all possible simple paths (i.e. paths with no cycles) from source $src_{i}$ to any safe node in ${\mathcal{S}}$.
@@
Proof of Lemma 1 is provided in the Appendix. Let $a_{-i}$ denote actions of all players excluding player $i$ in outcome $a$, i.e. $a_{-i}=(a_{1},...,a_{i-1},a_{i+1},...,a_{n})$. Then:
@@
To quantify the quality of equilibria, Price of Anarchy (PoA) and Price of Stability (PoS) is defined as follows:
@@
Figure 1(a) (evacuation network) and 1(b) (utility table) show an example instance of egres. Here, we have two players: $P_{0}$ and $P_{1}$, corresponding to the two source nodes 00 and $1$ in Figure 1(a). There is only one safe node ($A$) in this instance. Each player has $1$ evacuee under her. Figure 1(b) shows the possible actions of the two players. Actions of $P_{0}$ are shown in rows, and actions of $P_{1}$ in columns. For instance, the action $[0,2,A],[(0,1)]$ of $P_{0}$ means: $P_{0}$ has chosen the route $[0,2,A]$, and one evacuee leaves from node 00 at timestep 00. Each cell of the utility table in Figure 1(b) corresponds to an outcome. The first and second value in each cell are the utility values of player $P_{0},P_{1}$ respectively. These values are calculated according to Equation 1.
@@ IV-A Existence of Pure Strategy Nash Equilibrium
Corollary 1 indicates that an equilibrium can be as good as an optimal outcome. Corollary 2, on the other hand, has important implications about the computation of equilibria. We discuss this further in Section IV-C.
@@ IV-C1 Calculating Best Response
1 $\mathcal{F}\leftarrow$ Directed forest constructed using routes in $a_{-i}$.
@@ V-A Quality of the Equilibria
Figure 7: Box-plot showing the ratio of cost at equilibrium to a lower bound of optimal cost, of the $30$ equilibrium outcomes calculated by saa, for our two problem instances. The closer the ratio value to $1$, the better.
@@
Since, saa provides us equilibria that have confluent routes, we use the optimal confluent routes and schedule as our baseline. We use the Mixed Integer Program (MIP) formulation provided by Islam and Chen et al. [4] to find the optimal confluent routes and schedule for our two problem instances. We solved the MIPs with the MIP solver Gurobi [23]. However, since finding an optimal outcome is NP-hard and also hard to approximate [4], the solver was not able to provide us with exact optimal solutions. The solver did provide us with lower-bounds on the optimal cost (i.e. lower-bound on $\min_{a\in A}C(a)$). We denote this lower-bound by $opt_{lb}$, i.e. $opt_{lb}\leq\min_{a\in A}C(a)$. Let the cost (i.e. sum of the evacuation time of all evacuees) in an equilibrium outcome be denoted by $obj_{eq}$. We then look at the ratio $obj_{eq}/opt_{lb}$ to quantify the quality of the equilibria found by saa. For example, if the ratio is $2$ for an equilibrium outcome, then it means the cost at that equilibrium is at-most twice the cost at an optimal solution. The closer the ratio value to $1$, the better.
@@ V-B Scalability of saa
MIP solve time
$28$ $>120$ $>120$
SAA avg. run
time (minutes)
$14$ $20$ $27$
@@
MIP solve time
$56$ $>120$ $>120$
SAA avg. run
time (minutes)
$34$ $36$ $37$
@@ VI Conclusion
Acknowledgment
This work was partially supported by University of Virginia Strategic Investment Fund SIF160, the NSF Grants: CCF-1918656, OAC-1916805, RISE-2053013, and the NASA Grant 80NSSC22K1048.